You need to prepare a 47 mL of a 24% syrup/soda "prescription" . To do this, you can use the 80% syrup solution and the 10% soda solution that you have in stick. How many mL of each solution do you need to create this final solution?
Answer by fractalier(6550)

Call the amount of 80% solution you use, x.
Then 47-x is the amount of 10% solution used.
The set up is then
.80x + .10(47-x) = .24(47)
.80x + 4.7 - .10x = 11.28
.70x = 6.58
x = 9.4 mL of 80%
47-x = 37.6 mL of 10%